Define moment ofinertia.explain the physical significance of it.

Grade:12

2 Answers

Arun
5 years ago
moment of inertia is a measure of how difficult it is to rotate a particular body about a given axis. significance: Greater the mass concentrated away from the axis, greater the moment of inertia.
Rajdeep
5 years ago
HELLO THERE!
 
Well, your question is too long to answer, as Moment of Inertia is a complete topic, but here I’m explaining you what is Moment of Inertia with an example:
 
Moment of inertia, also called angular mass or rotational inertia can be defined with respect to a rotation axis, as a quantity that decides the amount of torque required for a desired angular acceleration or a property of a body due to which it resists angular acceleration. The formula for moment of inertia is the “sum of the product of mass” of each particle with the “square of its distance from the axis of the rotation”. The formula of Moment of Inertia is expressed as I = Σ miri2.
 
Consider an example here:
When you board a moving train, you experience a force that pushes you backward. That is because before boarding the train you were at rest. As soon as you board the moving train, your lower body comes in contact with the train but your upper body is still at rest. Therefore, it gets pushed backward, that is, it resists change in its state.
 
You need to learn the Moment of Inertia of some specific rigid bodies. I’m giving you some of them:
 
Cylinder = MR2/2
Solid sphere = 2/5 MR2
Hollow sphere = 2/3 MR2
Solid cone = 3/10 MR2
Hollow cone =MR2/2
 
HOPE MY ANSWER IS SATISFACTORY...
Thanks!